Circular No. 6371

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 J. Kommers and W. H. G. Lewin, Massachusetts Institute of Technol-

ogy; C. Kouveliotou and M. H. Finger, Universities Space Research Association; G. J. Fishman, C. A. Wilson, and R. B. Wilson, Marshall Space Flight Center; M. S. Briggs, University of Alabama in Huntsville (UAH); and J. van Paradijs, University of Amsterdam and UAH, report on the 2.1-Hz pulsations: “For a sample of 61 bursts from GRO J1744–28 observed by BATSE over a 3-month period (Jan.–Mar.), the average value of the pulsed fraction within bursts in the range 20–60 keV is 14 ± 1 percent (rms variation). The average value for the persistent emission observed by BATSE (covering the period Jan. 10–16) is 16.8 ± 0.3 percent (rms) in the energy range 20-40 keV. By comparison, the pulse fraction in the persistent emission observed by the PCA instrument (M. J. Stark, private communication) onboard XTE on Feb. 19 is 8 percent (rms) in the range 2–12.5 keV and 12 percent (rms) for E > 12.5 keV. No background subtraction or correction for the PCA response was performed. The pulse fraction in the upper energy range is therefore an underestimate. Dotani et al. (ASCA) report a ‘peak-to-peak’ pulsed fraction in the persistent emission of 10 percent in the band 2–10 keV (IAUC 6337, and T. Dotani, private communication). For comparison with the numbers given above, we note that this converts to a pulse fraction of roughly 3.5 percent (rms). [For an approximately sinusoidal pulse shape, as is the case here, the peak-to-peak pulse fraction is 2(2)^{0.5} times the rms pulse fraction.] From the combined BATSE-XTE-ASCA observations, it is very clear that there is a strong energy dependence of the pulsed fraction from GRO J1744–28.”

COMET C/1996 B2 (HYAKUTAKE) S. Ganesh, A. Chitre, H. O. Vats, K. S. Baliyan, U. C. Joshi, M. R.

Deshpande, N. M. Vadher, and A. B. Shah, Physical Research Laboratory (PRL), report: “Photopolarimetric observations of comet C/1996 B2 were made with the PRL 1.2-m telescope (+ optical polarimeter) at Gurushikhar, through IHW continuum filters (u, 365–368 nm; b, 484.5–486.5 nm; r, 684– 689 nm), with an aperture of 31′′ centered on the nucleus. On Mar. 13.97 UT, the polarization percentage (estimated error) and polarization angle for the filters mentioned above were, respectively: 6.38 (0.50), 15^{o}; 5.78 (0.22), 14^{o}; 6.32 (0.29), 14^{o}. On Mar. 18.88, the values were, respectively: 5.04 (0.24), 3^{o}; 5.30 (0.14), 6^{o}; 6.00 (0.22), 6^{o}.”

1996 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green